The Target Copy Company is contemplating the replacement of its old printing machine with a new model costing $ 7 0,000. The old machine, which originally cost $40,000, has 6 years of expected life remaining and a current book value of $ 19 ,000 versus a current market value of $ 22 ,000. Target's corporate tax rate is 40 percent. If Target sells the old machine at market value, what is the initial after-tax outlay for the new printing machine? Cash outflow must be a negative number! Round it a whole dollar and do not include the $ sign.
Answer #1

INITIAL AFTER TAX OUTLAY FOR NEW MACHINE = COST OF NEW MACHINE + AFTER TAX SALVAGE OLD MACHINE

THE BOOK VALUE OF OLD MACHINE = 19000

THE MARKET VALUE OF OLD MACHINE = 22000

SO THERE IS A PROFIT OF 22000 - 19000 = 3000 ON SALE OF OLD MACHINE

TAX PAYABLE ON PROFIT ON SALE OF OLD MACHINE = 40% X 3000 = 1200

SO AFTER TAX SALVAGE VALUE = MARKET VALUE - TAX PAYABLE ON PROFIT ON SALE

AFTER TAX SALVAGE VALUE = 22000 - 1200 = 20800

INITIAL AFTER TAX OUTLAY FOR NEW MACHINE = -70000 + 20800 = -49200

Answer : - 49200 (Thumbs up please)